# Coherent Control of an Optical Quantum Dot Using Phonons and Photons

### v1

@@ Abstract
Genuine quantum-mechanical effects are readily observable in modern optomechanical systems comprising bosonic (“classical”) optical resonators. Such systems have enabled laser-cooling of mesoscopic objects, generation of squeezed light, and the conversion of photons between microwave and optical modes. Here we demonstrate unique advantages of optical two-level systems, or qubits, for optomechanics. The qubit state can be coherently controlled using an immense variety of driving schemes including both phonons and resonant or detuned photons. We experimentally demonstrate this using charge-controlled InAs quantum dots (QDs) in surface-acoustic-wave resonators. Time-correlated single-photon counting measurements reveal the control of QD population dynamics using engineered optical pulses and mechanical motion. As a first example, we show how this can improve signal-to-background scattering in microwave-to-optical transduction processes. Specifically, we tailor the scheme so that mechanically assisted photon scattering is enhanced over the direct detuned photon scattering from the optical system. These differences are greatly amplified by strategic temporal pulse shaping. Quantum-mechanical calculations show good agreement with our experimental results and provide guidance for adapting these schemes to small phonon occupancies relevant for microwave-to-optical quantum transduction.
@@
Optomechanics underlies major technological advancements ranging from single particle and atom trapping [1] to gravitational-wave detection [2, 3]. As detection sensitivities and fabrication capabilities have improved, cavity optomechanics has been extended to explore and exploit more subtle quantum-mechanical effects [2, 4], enabling laser-cooling of mesoscopic objects [5, 6, 7, 8], new sources of squeezed light [9], transduction between microwaves and optical light [10, 11, 12, 13, 14, 7, 15], and entanglement between optical and mechanical [16, 17] or optical and microwave modes [18, 19]. Notably, all such previous demonstrations have exploited bosonic (“classical”) optical resonators, for example, whispering-gallery-mode resonators [5], Fabry-Perot resonators [10, 8], and photonic-crystal nanobeams [6, 11, 12, 13, 19, 15]; their purpose is essentially to confine the optical field to the region where the optomechanical interaction occurs, or to enforce frequency-dependent optical admittances that depend sensitively on mechanical motion. For applications where single-quantum interactions are most relevant, such as quantum microwave-optical transduction, single-phonon generation [20] or single-photon generation [21], optical two-level systems (TLSs) — appropriately referred to as qubits — may be beneficial because they inherently limit optical interactions to single-photon levels. They also enable coherent control schemes that are not possible in bosonic optical resonators. Optical TLSs and related three-level systems can resonantly couple with mechanical modes, offering potential routes to efficient optomechanical cooling and phonon “lasing” [22].
@@
Semiconductor quantum dots (QDs) are excellent realizations of optical TLSs [23, 24], exhibiting nearly gigahertz emission rates, near-unity quantum yields, near-transform-limited linewidths, high single-photon purity and high single-photon indistinguishability [25]. They have been used to generate nonclassical states of light [26], spin-photon entanglement [27, 28, 29], and photon-photon entanglement [30, 31]. Demonstrations of unconventional coherent excitation schemes using two ultrafast red-detuned laser pulses highlight the remarkable degree of control possible with QDs [21, 32]. Recent efforts involve integrating QDs into larger multifunctional hybrid quantum systems [33, 34, 35, 36]. QDs and other solid-state quantum emitters, e.g., diamond color centers, have more recently been explored as light-matter intermediaries in quantum optomechanical systems [22, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. Oscillations between excited orbital states in diamond color centers have been observed using multi-phonon processes and introduced as a method for dynamically mitigating decoherence [49]. To date, we know of no demonstrations of coherent optical and mechanical control of a coupled QD-phonon system. Such an effect could be used to improve signal-to-background performance in applications such as microwave-to-optical transduction by controlling the probabilities of direct-excitation scattering compared to phonon-assisted scattering from the QD. These same methods may be immediately applied to “swing-up” schemes to coherently prepare the QD’s exciton (biexciton) state for single-photon (entangled-photon) generation using a single nonresonant pump which can then be spectrally filtered [50].
Here, we use frequency- and time-tailored optical pulses, driven surface acoustic wave (SAW) cavity phonons, and time-domain optical measurements to experimentally demonstrate phonon-assisted coherent dynamics in a single QD. That is, QD excitation dynamics are controlled using both phonons and photons. Mechanical motion has large effects on TLS trajectories up to approximately 1 ns after the onset of optical driving, even after substantial radiative relaxation. As a first example — relevant to microwave-to-optical transduction — we show that optical pulse shapes and durations can be optimized such that phonon-assisted scattering is substantially enhanced compared to background scattering from direct detuned excitation. Quantum-mechanical calculations show good agreement with experiment, together indicating that large signal-to-background improvements are possible in these systems with optimized pulse shaping and device performance. There is no analogue in optomechanical systems comprising classical optical resonators.
I Visualizing phonon-assisted TLS dynamics
@@
Current realizations of hybrid QD-mechanical systems are operated at timescales much longer than relevant dephasing times, i.e. via continuous-wave (CW) pumping [41, 42, 34, 43, 47], or with nonresonant optical pumping [51, 38, 44], and therefore these coherent properties are obscured. In a few cases, unique TLS signatures have been observed through second-order photon correlation measurements [41, 47]. Dynamics in phonon-assisted scattering have been observed, but the optical TLS coherence was not exploited or explored [42]. In the CW case, the QD merely plays a role analogous to the optical resonators in the above-mentioned bosonic systems. In fact, the two-level nature of the QD could even be considered a disadvantage in these cases, as saturation can occur under the strong optical pumping required to get significant optomechanical coupling rates. This is remedied by using short optical pulses and well-known coherent control methods, as we show here.
Our system comprises a single optical QD with transition frequency $\omega_{\text{QD}}$ subject to an optical drive (frequency $\omega_{\text{pump}}$=$\omega_{\text{QD}}$+$\Delta$ with detuning $\Delta$) in a single-mode SAW cavity (frequency $\omega_{\text{SAW}}$). The SAW strain field parametrically modulates the QD’s frequency via deformation potential coupling [52, 37, 53, 54]. This coupling is described by a rate $g$, leading to an instantaneous QD frequency $\omega_{\text{QD}}$+$g\cos(\omega_{\text{SAW}}t)$. Importantly, this phonon-induced frequency modulation leads to discernible variations in the TLS occupancy when the QD is optically driven (Fig. 1). In practice, an optimized pulse shape would be developed to prepare the TLS in a desired state, then successive processes would be triggered on photons collected after this pulse. A more detailed theoretical description of this system is provided in Methods and in the Supplementary Information.
@@
Figure 1: (a) Top row: Bloch-sphere representation of the two-level system time evolution for: $i$) direct resonant excitation ($\Delta$=0), $ii$) direct red-detuned excitation ($\Delta$=$-\omega_{\text{SAW}}$), $iii$) phonon-assisted red-detuned excitation assuming a large coherent phonon population ($\Delta$=$-\omega_{\text{SAW}}$, $g/2\pi$=2 GHz). In all cases a constant $\Omega_{0}/2\pi$=1.5 GHz is used with the TLS initiated in its ground state at time $t$=0. Bottom row: Expected excited-state occupancy $\langle\rho_{ee}(t)\rangle$ for the same three cases. Calculations in panel a) do not include radiative relaxation or dephasing. (b) Time-dependent comparison of $\langle\rho_{ee}(t)\rangle$ between cases $ii$) (brown dashed curve) and $iii$) (red). Case $iii$) is shown for $g/2\pi$=2 GHz (pink) and $g/2\pi$=100 MHz (red) corresponding approximately to approximately $10^{3}$ phonons with $g_{0}/2\pi$=3 MHz. This calculation assumes an idealized square optical pulse shape and a radiative relaxation rate of $2\pi\times$320 MHz with no additional dephasing.
II Characterizing basic device performance
A fabricated QD-SAW optomechanical device is shown in Fig. 2a. Complete fabrication details are described in ref [46]. SAWs are confined between two acoustic mirrors forming an acoustic Fabry-Perot cavity at frequency $\omega_{\text{SAW}}/2\pi$$\approx$3.5 GHz with a quality factor $Q$$\approx$20,000 and finesse $\approx$130. The cavity both enhances the QD-SAW interaction compared to that on a bare surface [43] and allows us to parameterize the QD-SAW interaction by the single-phonon coupling rate $g_{0}$ and equilibrium phonon number $\bar{n}_{\text{SAW}}$. The SAW cavity mode is coherently populated with $\bar{n}_{\text{SAW}}$ phonons by driving an interdigital transducer (IDT) with an external microwave source.
@@ II Characterizing basic device performance
Figure 2: (a) Optical microscope image of the device. IDT: interdigital transducer. CPW: coplanar waveguide. (b) Schematic cross-section (not to scale) of the region under the Schottky top gate (indicated by a dotted white line in panel a). A single InAs QD mediates an optomechanical interaction between SAWs ($\omega_{\text{SAW}}$) and the optical pump ($\omega_{\text{pump}}$). DBR: distributed Bragg reflector. (c) Frequency-domain illustration of the system with a pulsed pump bandwidth of 600 MHz (black curve, resonant; red curve, red detuned) and a QD linewidth of 450 MHz (gray region). Three different scattering processes are illustrated with rates $\Omega_{0}$ (resonant direct excitation), $\Omega$ (detuned direct excitation) and $\Gamma$ (detuned phonon-assisted excitation), described in the main text. (d) Two types of resonance fluorescence (‘RF’) measurements from a single QD. Top panel: Total scattered photon counts while slowly varying the CW pump detuning $\Delta$ without (brown) and with large steady-state phonon occupancies ($g/2\pi$=1.09 GHz, dark red; $g/2\pi$=1.93 GHz, lighter red). Bottom panel: Frequency spectrum of scattered photons under red-detuned ($\Delta$=$-\omega_{\text{SAW}}$) CW optical pumping. The $g$ values and corresponding colors are the same in both panels. In each panel, spectra have been normalized to the maximum counts in the dataset. Gray region: pump background. (e) Exemplary experimental time-domain optical signal (counts per time bin) under $\Delta$=0 pulsed excitation acquired with a 15-second integration time. Important experimental features are illustrated. Blue curve: QD luminescence signal with decay time $\tau_{\text{QD}}$=1 ns. Gray region; pump background, showing the pulse shape and duration $\tau_{\text{pulse}}$$\approx$800 ps.
For good QD performance, QDs are electrostatically stabilized and controlled by a vertical diode (Fig. 2b) [55, 46]. This enables tuning of the QD’s charge state and wavelength with applied gate bias voltage (Supplementary Information) [56, 57] and reduces QD absorption linewidths to $\gamma_{\text{QD}}/2\pi$$<$350 MHz in best cases [58, 59, 57]. Under CW (narrowband) pumping, direct and phonon-assisted excitation can be fairly well isolated by changing the pump frequency. Under pulsed pumping, the increased pump bandwidth makes it more difficult to isolate these processes spectrally (Fig. 2c), but it opens the new possibility of coherent population control through pulse tuning as we next demonstrate.
@@
We interrogate a singly charged QD at 975.934 nm with an approximately 800 MHz linewidth (full width at half maximum) measured over a timescale of several seconds. This broadening beyond the radiative lifetime limit is likely due to slow (sub-MHz) spectral wandering [60, 59], as we will describe in more detail in a later section. Nonetheless, our system satisfies the condition $\omega_{\text{SAW}}$$>$$\gamma_{\text{QD}}$ — known as the resolved-sideband regime — as verified by measuring the absorption (Fig. 2d; top panel) and scattering spectra (Fig. 2d; bottom panel) of the single QD under SAW modulation. Well-defined spectral sidebands at $\Delta$=$\pm$$\omega_{\text{SAW}}$ correspond to single-phonon-assisted optical scattering from the QD [52, 61, 43]. Our main optical measurements involve exciting the QD with frequency-tuned and time-tailored optical pulses (Methods). A CW pump with variable detuning $\Delta$ originates from a narrow-linewidth ($\sim$10 kHz) laser. Pulses with durations of approximately 1 ns are generated by passing the CW pump through two electro-optic modulators and a tunable spectral filter (Methods). The time distribution of scattered photons is resolved using time-correlated single-photon counting techniques, as illustrated in Fig. 2e.
III Resolving phonon-assisted excitation dynamics
@@
The long ($\approx$1 ns) pulse duration allows us to directly observe, within approximately 10 seconds, the QD’s average population dynamics as the QD emits photons at random times during each pulsed-excitation event. These time-domain traces of the QD’s dynamics are each composed of several million photon detection events (Methods). Series of such measurements under red-detuned ($\Delta$=$-\omega_{\text{SAW}}$) pumping with $g$ varying between approximately 0 (only thermal occupancy) and 2 GHz are shown in Figs. 3a-b. Measured count rates are converted to estimated QD occupancies using a calibration procedure described in Supplementary Information. For the remainder of this article, we will refer to the ratio of occupancies as the “enhancement”, $c_{g}(t)$=$s_{g}(t)/s_{g=0}(t)$ where $s_{g}(t)$ is the signal at time $t$ with coupling rate $g$. This enhancement measures how well the presence or absence of phonons can be detected with the optical signal.
@@
Figure 3: (a,b) Experimental QD dynamics under red-detuned ($\Delta$=$-\omega_{\text{SAW}}$) pumping (cases $ii$ and $iii$) with two different optical pulse shapes: a) a pulse with a gradually decreasing intensity; b) a nearly square pulse with an abruptly decreasing intensity. Top row: measured pulse intensities for the two cases. Approximate peak resonant Rabi frequencies are specified in each panel. Center row: measured time-resolved QD occupancy. Trajectories for a range of QD-SAW coupling rates $g$ are shown ($g$ specified next to a subset of the curves). Bottom row: occupancy ratios (“enhancements”), $c_{g}(t)$, for each of the $g$$>$0 curves compared to the $g$=0 curve. Gray region indicates where the dynamics are dominated by free evolution (decay) of the TLS. (c,d) Numerical calculations corresponding approximately to the two experimental scenarios presented in panels a,b. Calculations assume radiative relaxation at 320 MHz and no additional dephasing. Calculations for the “exponential” pulse are truncated at 1.25 ns since oscillations would formally continue indefinitely leading to indefinitely large enhancements at long times. Note that the occupancy ratio in panel c is presented on two independently scaled axes because maxima span several orders of magnitude.
@@
We first discuss results using a pulse with a gradually decreasing intensity (Fig. 3a). Experimentally measured pulse shapes are shown in the top panels. The QD occupancy (Fig. 3a: center panel) with no added phonons (brown curve) shows several Rabi oscillations at frequency $\Omega$, which is dominated by the detuning. Obvious changes in the dynamics are resolved when the SAW cavity is occupied such that $g/2\pi$=$770$ MHz (lowest red curve). These differences become more pronounced as $g$ increases, where the QD occupancy is estimated to reach 0.075 for $g/2\pi$=$2.4$ GHz at 800 ps. The data reveal that there are several pulse durations where phonon-assisted excitation (red curves) dominates substantially over the bare QD excitation (brown curve) — namely around 300 ps and 600 ps — largely governed by times when the bare QD occupancy swings down to approximately zero. These measured enhancements $c_{g}(t)$ reach between 12 and 45 at 600 ps, depending on the SAW cavity occupancy (Fig. 3a; bottom panel). To illustrate the importance of the pulse shape on the measured enhancements, we perform a similar set of measurements using an (approximately) square optical pulse (Fig. 3b). Although obvious changes in the dynamics are again observed with $g/2\pi$ as small as 340 MHz (bottom panel; red curve), the measured enhancements are considerably smaller than in the “exponential” pulse case, remaining near unity in all cases presented here.
@@
The benefit of a gradually decreasing pulse intensity — as compared to a square pulse for example — becomes clear upon inspection of the bare QD dynamics. Numerical calculations of QD dynamics are presented in Figs. 3c-d, roughly corresponding to the two experimental cases shown in Figs. 3a-b (Methods). For a pulse with constant intensity (Fig. 3d; brown curve), the bare QD occupancy oscillates around and approaches a constant central value as radiative relaxation (or dephasing in general) proceeds. Consequently, successive minima in the bare QD’s occupancy grow increasingly large as the pulse continues. This behavior is also observed in our experiments (Fig. 3b). Enhancements remain smaller than 10 throughout the pulse duration. If the pulse intensity is instead gradually decreased soon after the onset, minima in the bare QD’s occupancy can grow increasingly close to zero over the pulse duration (Fig. 3c; brown curve). Enhancements thus increase over the course of the QD’s lifetime or longer, approaching maximum values near 1000 for $g/2\pi$$\gtrsim$1 GHz at long times. Calculations presented in Fig. 3 assume radiative relaxation and no additional dephasing (Methods); the effect of additional dephasing is discussed in the next section.

### v2

@@ Abstract
Genuine quantum-mechanical effects are readily observable in modern optomechanical systems comprising bosonic (“classical”) optical resonators. Here we describe unique features and advantages of optical two-level systems, or qubits, for optomechanics. The qubit state can be coherently controlled using both phonons and resonant or detuned photons. We experimentally demonstrate this using charge-controlled InAs quantum dots (QDs) in surface-acoustic-wave resonators. Time-correlated single-photon counting measurements reveal the control of QD population dynamics using engineered optical pulses and mechanical motion. As a first example, at moderate acoustic drive strengths, we demonstrate the potential of this technique to maximize fidelity in quantum microwave-to-optical transduction. Specifically, we tailor the scheme so that mechanically assisted photon scattering is enhanced over the direct detuned photon scattering from the QD. Spectral analysis reveals distinct scattering channels related to Rayleigh scattering and luminescence in our pulsed excitation measurements which lead to time-dependent scattering spectra. Quantum-mechanical calculations show good agreement with our experimental results, together providing a comprehensive description of excitation, scattering and emission in a coupled QD-phonon optomechanical system.
@@
Optomechanics underlies major technological advancements ranging from single particle and atom trapping [1] to gravitational-wave detection [2, 3]. As detection sensitivities and fabrication capabilities have improved, cavity optomechanics has been extended to explore and exploit more subtle quantum-mechanical effects [2, 4], enabling laser-cooling of mesoscopic objects [5, 6, 7, 8], new sources of squeezed light [9], transduction between microwaves and optical light [10, 11, 12, 13, 7, 14], and entanglement between optical and mechanical [15, 16] or optical and microwave modes [17, 18]. Notably, all such previous demonstrations have exploited bosonic (“classical”) optical resonators, for example, whispering-gallery-mode resonators [5], Fabry-Perot resonators [10, 8], and photonic-crystal nanobeams [6, 11, 12, 13, 18, 14]; their purpose is essentially to confine the optical field to the region where the optomechanical interaction occurs, or to enforce frequency-dependent optical admittances that depend sensitively on mechanical motion. Genuine quantum control of mechanical motion, however, requires an external two-level system (TLS) for state preparation and readout. This has in fact been realized in microwave electromechanics by coupling superconducting qubits to mechanical modes, sparking the field of quantum acoustodynamics [19, 20]. Analogous interactions have not yet been established between mechanical modes and optical TLSs, although this would be beneficial in applications such as quantum microwave-optical transduction, single-phonon generation [21] or single-photon generation [22]. Optical TLSs inherently limit optical interactions to single-photon levels and enable coherent control protocols that are not possible in bosonic optical resonators. Optical TLSs and related three-level systems can resonantly couple with mechanical modes, offering potential routes to efficient optomechanical cooling and phonon “lasing” [23].
@@
Semiconductor quantum dots (QDs) are excellent realizations of optical TLSs [24, 25], exhibiting nearly gigahertz emission rates, near-unity quantum yields, near-transform-limited linewidths, high single-photon purity and high single-photon indistinguishability [26]. They have been used to generate nonclassical states of light [27], spin-photon entanglement [28, 29, 30], and photon-photon entanglement [31, 32]. Demonstrations of unconventional coherent excitation schemes using two ultrafast red-detuned laser pulses highlight the remarkable degree of control possible with QDs [22, 33]. Recent efforts involve integrating QDs into larger multifunctional hybrid quantum systems [34, 35, 36, 37]. QDs and other solid-state quantum emitters, e.g., diamond color centers, have more recently been explored as light-matter intermediaries in quantum optomechanical systems [23, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. Oscillations between excited orbital states in diamond color centers have been observed using multi-phonon processes and introduced as a method for dynamically mitigating decoherence [50]. To date, we know of no demonstrations of coherent optical and mechanical control of a coupled QD-phonon system.
Here, QD excitation dynamics are controlled using both phonons and photons. Mechanical motion has large effects on Bloch-sphere trajectories up to approximately 1 ns after the onset of optical driving, even after substantial radiative relaxation. As a first example, we establish a protocol directly applicable to optically heralded entanglement of remote acoustic resonators. In such a protocol, detection of an optomechanically scattered photon heralds the successful preparation of the mechanical state, but zero-phonon scattering events are “false alarms” that degrade state preparation fidelity. We show that optical pulse shapes can be optimized such that phonon-assisted scattering is substantially enhanced compared to background scattering from direct detuned excitation, mitigating the deleterious effect of such background photons. There is no analogue in optomechanical systems comprising classical optical resonators. Finally, we analyze the frequency spectrum of photons from a QD prepared using either direct or phonon-assisted excitation. These spectra reveal distinct Rayleigh-like scattering channels and luminescence channels which in the pulsed-excitation case lead to time-dependent spectra and verify successful optomechanical excitation of the QD. These results provide a comprehensive picture of excitation, scattering, and emission processes in a coupled phonon-QD optomechanical system.
I Understanding phonon-assisted TLS dynamics
@@
Current realizations of hybrid QD-mechanical systems are operated at timescales much longer than relevant dephasing times, i.e. via continuous-wave (CW) pumping [42, 43, 35, 44, 48], or with nonresonant optical pumping [51, 39, 45], and therefore these coherent properties are obscured. In a few cases, unique TLS signatures have been observed through second-order photon correlation measurements [42, 48]. Dynamics in phonon-assisted scattering have been observed, but the optical TLS coherence was not exploited or explored [43]. In the CW case, the QD merely plays a role analogous to the optical resonators in the above-mentioned bosonic systems. The two-level nature of the QD could even be considered a disadvantage in these cases, as saturation can occur under the strong optical pumping required for significant optomechanical coupling rates. This is remedied by using short optical pulses and well-known coherent control methods, as we show here.
Our system comprises a single optical QD with transition frequency $\omega_{\text{QD}}$ subject to an optical drive of frequency $\omega_{\text{pump}}$=$\omega_{\text{QD}}$+$\Delta$ with detuning $\Delta$. It is coupled to phonons in a single-mode SAW cavity at frequency $\omega_{\text{SAW}}$ (Fig. 1a). The SAW strain field parametrically modulates the QD frequency via deformation potential coupling [52, 38, 53, 54]. Including this coupling, the QD-SAW system can be described by a ladder-like energy diagram (Fig. 1b). Diagonal transitions that add (remove) phonons to (from) the acoustic resonator are addressed with an optical drive at detuning $+\omega_{\text{SAW}}$ ($-\omega_{\text{SAW}}$) with a renormalized interaction rate of $\Gamma$ [2, 53]. Even with this detuning, however, there is a finite probability of exciting the QD “directly”, i.e., via vertical transitions involving no phonons. This process is inherent to the TLS and is described by the generalized Rabi rate $\Omega$=$\sqrt{\Omega_{0}^{2}+\Delta^{2}}$, where $\Omega_{0}$ is the resonant Rabi rate between the QD and pump field. This undesired excitation channel can be removed by controlling the system in the time domain. Fig. 1c presents simplified calculations illustrating the protocol. The zero-phonon (“direct”) excitation channel leads to occupancy oscillations of the $|e,n\rangle$ state at frequency $\Omega$ (Fig. 1c, black curve). Excitation via the phonon-assisted transition, which is in fact resonant with the pump, evolves differently (red curve). There are thus discrete pulse durations where the direct excitation channel is effectively eliminated, leaving only the phonon-assisted excitation. This phonon-assisted excitation process can alternatively be visualized on the bare QD’s Bloch sphere (Fig. 1d). Without phonon coupling, the QD’s state traces closed small circles near the south pole (ground state) of the Bloch sphere (black). The acoustic field perturbs this trajectory, generally increasing the QD’s occupancy while the optical drive remains active. The QD occupancy approaches unity given enough time and assuming sufficiently slow dephasing [55]. Further theoretical details are provided in the Supplementary Information.
Figure 1: (a) Illustration of the system. A QD is coupled to surface acoustic wave (SAWs) cavity phonons and is driven by an optical pump. (b) Schematic energy-level ladder diagram of the coupled QD and phonon resonator (frequency $\omega_{\text{SAW}}$). A pump laser with frequency detuning $\Delta$=$-\omega_{\text{SAW}}$ ($+\omega_{\text{SAW}}$) can address optical-frequency transitions by removing (adding) phonons to (from) the resonator. The coupling rate associated with these diagonal transitions is $\Gamma$=$g_{0}\Omega_{0}\sqrt{n}/\omega_{\text{SAW}}$, where $g_{0}$ is the single-phonon optomechanical coupling rate, $n$ is the cavity phonon number, and $\Omega_{0}$ ($\Omega$=$\sqrt{\Omega_{0}^{2}+\Delta^{2}}$) is the resonant (generalized) Rabi rate between the laser field and the QD. (c) Calculated excited-state occupancies of the states $|e,n\rangle$ (black) and $|e,n-1\rangle$ (red) from a square pulse initiated at time $t$=0 with the following system parameters: $\Delta$=$-\omega_{\text{SAW}}$=--$2\pi\times$3.5 GHz, $g_{0}\sqrt{n}$=$2\pi\times$1 GHz, $\Omega_{0}$=$2\pi\times$1 GHz. Vertical gray dashed lines indicate pulse durations where high-fidelity optomechanical state preparation is possible. (d) Bare QD Bloch-sphere trajectories for direct (black) and phonon-assisted (red) excitation process with $\Delta$=$-\omega_{\text{SAW}}$=--$2\pi\times$3.5 GHz, $g_{0}\sqrt{n}$=$2\pi\times$2 GHz, $\Omega_{0}$=$2\pi\times$1 GHz. Calculations in panels b) and c) assume no dephasing.
II Characterizing basic device performance
@@
Figure 2: (a) Optical microscope image of the device. IDT: interdigital transducer. CPW: coplanar waveguide. (b) Schematic cross-section (not to scale) of the region under the p-type top gate (indicated by a dotted white line in panel a). A single InAs QD mediates an optomechanical interaction between SAWs ($\omega_{\text{SAW}}$) and the optical pump ($\omega_{\text{pump}}$). DBR: distributed Bragg reflector. (c) Frequency-domain illustration of the system. Three different scattering processes are illustrated with rates $\Omega_{0}$ (resonant direct excitation), $\Omega$ (detuned direct excitation) and $\Gamma$ (detuned phonon-assisted excitation), described in the main text. (d) Two types of resonance fluorescence (‘RF’) measurements from a single QD. Left panel: Total scattered photon counts while slowly varying the CW pump detuning $\Delta$ without (brown) and with large steady-state phonon occupancies ($g/2\pi$=0.87 GHz, dark red; $g/2\pi$=1.55 GHz, lighter red). Right panel: Frequency spectrum of scattered photons under red-detuned ($\Delta$=$-\omega_{\text{SAW}}$=$-2\pi\times$3.58 GHz) CW optical pumping. The $g$ values and corresponding colors are the same in both panels. In each panel, spectra have been normalized to the maximum counts in the dataset. Gray regions: measured pump background.
@@
For good QD performance, QDs are electrostatically stabilized and controlled by a vertical p-i-n diode (Fig. 2b) [56, 47]. This enables tuning of the QD’s charge state and wavelength with applied gate bias voltage (Fig. S2) [57, 58] and reduces QD absorption linewidths to $\gamma_{\text{QD}}/2\pi$$<$350 MHz in best cases [59, 60, 58]. We interrogate a singly charged QD at 956.368 nm with an approximately 950 MHz linewidth (full width at half maximum) measured over a timescale of several seconds. This broadening beyond the radiative lifetime limit is likely due to slow (sub-MHz) spectral wandering [61, 60] rather than additional dephasing at the timescale of the QD’s lifetime (Supplementary Information). Nonetheless, our system satisfies the condition $\omega_{\text{SAW}}$$>$$\gamma_{\text{QD}}$, known as the resolved-sideband regime. Fig. 2c illustrates the spectral components and the three scattering processes (described in the previous section) of primary concern in our experiments. Operation in the resolved-sideband regime is verified with CW optical spectroscopies by measuring the absorption (Fig. 2d; left panel) and scattering spectra (Fig. 2d; right panel) of the single QD with and without SAW modulation. Well-defined spectral sidebands at multiples of $\omega_{\text{SAW}}$ correspond to single-phonon-assisted optical scattering from the QD [52, 62, 44]. To measure system dynamics, we excite the QD with frequency-tuned and time-tailored optical pulses (Methods). A CW pump with variable detuning $\Delta$ originates from a narrow-linewidth ($\sim$10 kHz) laser. Pulses with durations of approximately 100 ps to $>$1 ns are generated by passing the CW pump through two electro-optic modulators and a tunable spectral filter (Methods). The time distribution of scattered photons is resolved using time-correlated single-photon counting (TCSPC) techniques.
III Experimentally resolving phonon-assisted excitation dynamics
@@
Using near-resonant pulses with duration similar to the QD lifetime, we directly observe the QD’s average population dynamics as it emits photons at random times during each pulse. These time-domain traces of the QD’s dynamics are each composed of several million photon detection events (Methods). Series of such measurements under red-detuned ($\Delta$=$-\omega_{\text{SAW}}$) pumping with $g$ varying between approximately 0 (thermal phonon occupancy) and 1.55 GHz are shown in Figs. 3a-b. Measured count rates are converted to estimated QD occupancies using a calibration procedure (Fig. S3). For the remainder of this article, we will refer to the ratio of occupancies (minus one) as the “enhancement”, $c_{g}(t)$=$s_{g}(t)/s_{g=0}(t)-1$ where $s_{g}(t)$ is the photon signal at time $t$ with coupling rate $g$. This enhancement measures how well the presence or absence of phonons can be distinguished with the optical signal. If operated in the quantum regime, it is directly related to fidelity.
@@
Figure 3: (a,b) Experimental QD dynamics under red-detuned ($\Delta$=$-\omega_{\text{SAW}}$) pumping with two different optical pulse shapes: a) a pulse with a gradually decreasing intensity; b) a nearly square pulse with an abruptly decreasing intensity. Top row: measured pulse intensities for the two cases. Approximate peak resonant Rabi frequencies are specified in each panel. Center row: measured time-resolved QD occupancy. Trajectories for a range of QD-SAW coupling rates $g$ (different colors) are shown ($g/2\pi$ specified next to each curve in units of GHz). Bottom row: “enhancements”, $c_{g}(t)$, for each of the $g$$>$0 curves compared to the $g$=0 curve from the occupancy data shown in the center row. (c,d) Numerical calculations corresponding to the two experimental scenarios presented in panels a,b. Calculations assume radiative relaxation at $2\pi\times$320 MHz and small additional dephasing ($2\pi\times$60 MHz) to match the experiment. Fig. S5 shows the same data with a two-dimensional format.
@@
We first discuss measured dynamics using a pulse with a gradually decreasing intensity (Fig. 3a). Experimentally measured pulse shapes are shown in the top panels of Fig. 3a. The QD occupancy (Fig. 3a, center panel) with no added phonons ($g/2\pi$=0, brown curve) shows several Rabi oscillations at frequency $\Omega$, which is dominated by the detuning. Obvious changes in the dynamics are resolved when the SAW cavity is occupied such that $g/2\pi$=$490$ MHz. These differences become more pronounced as $g$ increases, where the QD occupancy is estimated to reach 0.035 for $g/2\pi$=$1.55$ GHz at 900 ps. The data reveal that there are several pulse durations where phonon-assisted excitation (red curves) dominates substantially over the bare QD excitation (brown curve) — namely around 300 ps, 550 ps, and 800 ps — largely governed by times when the bare QD occupancy swings down toward zero. These measured enhancements $c_{g}(t)$ (bottom panel) reach approximately 15 at 1.1 ns — and continue to increase smoothly as the pulse slowly wanes — but vary widely depending on the SAW cavity occupancy and pulse duration. To illustrate the importance of the pulse shape on the measured enhancements, we perform a similar set of measurements using an (approximately) square optical pulse (Fig. 3b). Although obvious changes in the dynamics are again observed when $g/2\pi$$>$0, the measured enhancements are considerably smaller than in the “decreasing pulse” case.
@@
The benefit of a gradually decreasing pulse intensity — as compared to a square pulse, for example — becomes clear upon inspection of the bare QD dynamics. Numerical calculations of QD dynamics are presented in Figs. 3c-d, corresponding to the two experimental cases shown in Figs. 3a-b (Methods). For a pulse with constant intensity (Fig. 3d), the bare QD occupancy oscillates around and approaches a constant central value (here, 0.06) as dephasing proceeds (brown curve). Consequently, successive minima in the bare QD’s occupancy grow increasingly large as the pulse continues. This behavior is well-resolved in our experiments (Fig. 3b). Enhancements thus remain small throughout the pulse duration. If the pulse intensity is instead gradually decreased soon after the onset, minima in the bare QD’s occupancy do not increase over the pulse duration (Fig. 3c; brown curve). Enhancements thus increase over the course of the QD’s lifetime or longer as the phonons have more time to interact with the optically driven QD.
There is a competition between the rate of decrease of the pump intensity and the QD dephasing. An ideal pulse would bring the QD occupancy down to zero before substantial dephasing. This can be accomplished by improving the QD coherence time, or by decreasing the pump intensity quicker than that demonstrated in Fig. 3a. Calculations presented in Fig. 3c,d assume radiative relaxation and small additional dephasing to match the experimental dynamics, although this additional dephasing likely imitates spectral wandering in this particular case. Reduced dephasing and no spectral wandering in an otherwise identical system substantially improves the enhancements (Fig. S6). With no dephasing nor spectral wandering, enhancements approach maximum values of approximately 1000 for $g/2\pi$$\gtrsim$1 GHz around 1.3 ns delay as the bare QD occupancy grows increasingly close to zero over the duration of the gradually decreasing pulse.